OpenStudy (angelwings996):

Use each recursive definition to write an explicit formula for the sequence. a_1=-5, a_n=a_(n-1)-1

OpenStudy (phi):

can you write down a few of the numbers in the sequence ? the first number is -5 what is the next number ?

OpenStudy (anonymous):

an = -4 + an-1

OpenStudy (angelwings996):

@cooki3 Is that the formula?

OpenStudy (anonymous):

YES

OpenStudy (angelwings996):

How did you get it?

OpenStudy (angelwings996):

@phi i don't know how to find the other numbers, do I plug in 2 for n to find the second number in the sequence and so one?

OpenStudy (anonymous):

a1=−2,a2=−4a1=−4a2∴an+1=−4an 4an = -4 + an-1

OpenStudy (angelwings996):

@cooki3 Thank you!! (:

OpenStudy (anonymous):

U WELCOME

OpenStudy (phi):

you have to learn how to interpret this \[ a_1=-5, \text{ }a_n=a_{n-1}-1 \] that says a_1 (the first term) is -5 the 2nd equation says: to get the next term, take the "current term" and subtract 1 if we replace n with 2, the 2nd equation says \[ a_2=a_{2-1}-1 \] or \[ a_2=a_{1}-1 \] you know a1 so you can figure out a2

OpenStudy (angelwings996):

Oh okay I see, thank you! @phi

OpenStudy (phi):

so what are the next few #'s in the sequence ?